Nancy Johnson invested $18,000. Part of it was invested at 3% annual simple interest, and the rest was invested at 4%. Her interest income for the first year was $650. How much did she invest at each rate?
This is what I tried: 3x + 4y = 650;
y=18x
3(18y) + 4y = 650
54y + 4y = 650 = 58y =650, y=11.206896
I got lost from there because when that would make X something not nice. Please help me?

Answer by rfer(10432)

.03x+.04(18000-x)=650
.03x+720-.04x=650
-.01x=-70
x=$7000 in 3%
18000-x=$11000 in 4%